# O. Ragniscoa, F.Zullob

### v1

@@ 5.1 An example for the discrete case
$\begin{split}&\epsilon_{n}=\frac{1}{\sqrt{2N}}\tilde{\epsilon}_{n},\;\\ &\eta_{n}=\sqrt{\frac{e^{2a}-1}{2(1-e^{-2aN})}}e^{-a|n|}\tilde{\epsilon}_{n}\\ &k_{n}=Ce^{-a|n|}\end{split}$ (5.1)
where $\tilde{\epsilon}_{n}$ is the sequence given in (3.20). Both $\epsilon_{n}$ and $\eta_{n}$ are normalized, in the $\ell^{2}$ norm, to the unity.
From the general formulae (3.30) and (3.34) we get
$\dot{Q}=P,\quad\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (5.2)
Let us look at the sum defining $\dot{P}$. By taking tha Taylor series of the exponential we can write
$\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\eta_{n}\frac{(Q\eta_{n}+\kappa_{n})^{k}}{k!}.$ (5.3)
@@
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\eta_{n}\frac{(Q\eta_{n}+\kappa_{n})^{k}}{k!}=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}$ (5.4)
@@
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{(1-\cos(\pi m))}{2}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}.$ (5.5)
We now recall, from (5.1), that when $m+1$ is even $\eta^{m+1}_{n}=w^{m+1}e^{-a(m+1)|n|}$, where $w=\sqrt{\frac{e^{2a}-1}{2}}$. By considering also also the explicit formula for $\kappa_{n}$ from (5.1) and by performing the sum over $n$ we can write
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}=\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}(1-\cos(\pi m))\frac{Q^{m}w^{m+1}_{n}C^{k-m}}{k!(e^{a(k+1)}-1)}$ (5.6)
@@
$\begin{split}&\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}(1-\cos(\pi m))\frac{Q^{m}w^{m+1}_{n}C^{k-m}}{k!(e^{a(k+1)}-1)}=\\ &=w\sum_{k}\frac{(C+Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}-w\sum_{k}\frac{(C-Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}\end{split}$ (5.7)
@@
$\dot{P}=1-w\sum_{k}\frac{(C+Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}+w\sum_{k}\frac{(C-Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}$ (5.8)
We notice that the series (5.8) can be also expressed in terms of the following function:
$F(z)=\sum_{n=1}r^{n}z^{r^{n}}=z\frac{d}{dz}\sum_{n=1}(z^{r^{n}}-1),$ (5.9)
@@
$\begin{split}&\dot{Q}=P,\\ &\dot{P}=1+w(F(e^{C-Qw})-F(e^{C+Qw}))\end{split}$ (5.10)
where the function $F(z)$ defined in (5.9) satisfies the functional equation
$rF(z^{r})=F(z)-rz^{r},\quad r\doteq e^{-a},$ (5.11)
@@
Alternatively but equivalently, the equations of motion (5.10) can be defined by the function $f(z)=F(\exp(z))$, given by the series
$f(z)=\sum_{k=0}\frac{z^{k}}{k!}\frac{r^{k+1}}{1-r^{k+1}}.$ (5.12)
@@
$\sum_{n=1}^{N}r^{n}\exp(zr^{n}),$ (5.13)
by expanding the exponential in its Taylor series, summing over $n$ and letting $N\to\infty$ one gets $F(\exp(z))=f(z)$, with $f(z)$ defined as in (5.12).
@@ 5.1.1 The numerics of the discrete case
$f(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\frac{1}{3^{\frac{k+1}{2}}-1}.$ (5.19)
@@
$|R_{N}|=|\sum_{k=N+1}^{\infty}\frac{z^{k}}{k!}\frac{1}{3^{\frac{k+1}{2}}-1}|\leq\frac{\sqrt{3}}{2}\sum_{k=N+1}^{\infty}\frac{(\frac{|z|}{\sqrt{3}})^{k}}{k!}=\frac{\sqrt{3}}{2}\frac{(\frac{|z|}{\sqrt{3}})^{N+1}}{(N+1)!}\exp(\frac{\theta|z|}{\sqrt{3}}),$ (5.20)
where $\theta\in(0,1)$ and we used the Lagrange form of the reminder for the expansion of $\exp(\frac{|z|}{\sqrt{3}})$. Just to get an idea, for $N=30$ the previous formula gives an $R_{30}$ of the order $10^{-8}$ for $z=10$ and $\theta=1$ and an $R_{40}$ of $10^{-16}$ for $N=40$ and the same values of $z$ and $\theta$, whereas the values of the sums from $k=0$ to $k=30$ and from $k=0$ to $k=40$ are about $2\cdot 10^{2}$. We use the series up to $N=40$ and integrate numerically the equations of motion with a standard Runge-Kutta-Fehlberg method with Maple. The Figure (2) shows the numerical integration of the system (5.18) with three different initial conditions: $(Q(0),P(0))=(10,0)$, $(Q(0),P(0))=(8,5)$ and $(Q(0),P(0))=(5,0)$. The corresponding values of the energies are $607.38102$, $207.09954$ and $33.39845$. The plots on the right, in blue, are the plots of the analytical curves corresponding to the level curves defined by the Hamiltonian in these three cases and coincide with the orbits. The curve in red is the curve defined by $\dot{P}=0$: the equilibrium position is where this curve intersect the $Q$ axis.
Figure 2: Plot of three orbits of the system (refnumeq): on the left there is the numeric, on the right the plot of the corresponding Hamiltonian level curves. In red the curve defined by $\dot{P}=0$.
@@
As specified in the Remark (3.8), the equations of motion (3.51) holds true also if the sequences $\eta_{n}$ and $\kappa_{n}$ belong to $\ell^{1}$. The properties of the orbits given in the Remark (3.6) or in the Remark (3.7) are also valid in these cases. As an example of the differential equations that one obtains, we can take $\epsilon_{n}$ like in (3.38) (hence $\kappa_{n}$ odd)
$\begin{split}&\eta_{n}=\frac{3\sigma}{\pi^{2}n^{2}},\;\;n\neq 0,\quad\eta_{0}=0,\quad\sigma^{2}=1\\ &\kappa_{n}=\frac{C\text{sgn}(n)}{n^{2}}\;\;n\neq 0,\quad\kappa_{0}=0.\end{split}$ (5.21)
From formula (3.51), to get $\dot{P}$ we need to evaluate the following sum
$\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}),$ (5.22)
and, by using explicitly the relations (5.21) we get
$\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\frac{3\sigma}{\pi^{2}}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\left(\exp\left(\frac{\frac{3\sigma}{\pi^{2}}Q+C}{n^{2}}\right)+\exp\left(\frac{\frac{3\sigma}{\pi^{2}}Q-C}{n^{2}}\right)\right).$ (5.23)
@@
$\zeta_{even}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\zeta(2k+2),$ (5.24)
@@
$\dot{Q}=P,\quad\dot{P}=1-\frac{3\sigma}{\pi^{2}}\zeta_{even}\left(\frac{\frac{3\sigma}{\pi^{2}}Q+C}{n^{2}}\right)-\frac{3\sigma}{\pi^{2}}\zeta_{even}\left(\frac{\frac{3\sigma}{\pi^{2}}Q-C}{n^{2}}\right)$ (5.25)
@@ 5.3 An example for the continuous case
$\begin{split}&\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\tilde{\epsilon}^{2}=1.\\ &\eta(x)=\frac{a}{\sqrt{1-\exp(-2a^{2}L)}}\exp(-a^{2}|x|)\tilde{\epsilon}(x),\\ &\kappa(x)=C\exp(-a^{2}|x|).\end{split}$ (5.26)
Let us look at the equations of motion for $P$ (4.4) and $Q$ (4.5). From the general formulae (4.36) and (4.41) we get
$\dot{Q}=P,\quad\dot{P}=1-\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi$ (5.27)
@@
$\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi=\sum_{n=0}^{\infty}\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}$ (5.28)
Since $\eta$ is an odd function, whereas $\kappa$ is even, only when $j$ is odd there is a contribution different from zero. So, by taking into account the expressions (5.26) and by noticing that $\tilde{\epsilon}^{j+1}=1$ for $j$ odd we can write:
$\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}=\sum_{j=0}^{n}{n\choose j}\frac{1-\cos(\pi j)}{2}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}C^{n-j}\alpha^{j+1}e^{-(n+1)\alpha^{2}|\xi|}}{n!}$ (5.29)
@@
$\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}=\sum_{j=0}^{n}{n\choose j}\frac{\alpha^{j-1}Q^{j}C^{n-j}(1-\cos(\pi j))}{(n+1)!}.$ (5.30)
The sum over $j$ can be carried out by using the formula for the binomial expansion. The sum in equation (5.30) is just given by
$\frac{(C+Qa)^{n}}{a(n+1)!}-\frac{(C-Qa)^{n}}{a(n+1)!}.$ (5.31)
Going back to equation (5.28) we can write
$\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi=\sum_{n=1}\left(\frac{(C+Qa)^{n}}{a(n+1)!}-\frac{(C-Qa)^{n}}{a(n+1)!}\right),$ (5.32)
@@
$\dot{Q}=P,\quad\dot{P}=1-\frac{e^{C+aQ}-1}{a(C+aQ)}+\frac{e^{C-aQ}-1}{a(C-aQ)}.$ (5.34)
@@
$\frac{e^{C+aQ}-1}{a(C+aQ)}+\frac{e^{C-aQ}-1}{a(C-aQ)}=1.$ (5.35)
As expected, this equation has always just one real root. For the sake of simplicity, let us just set $a=1$. Then, both the functions $\frac{e^{x}-1}{x}$ and $\frac{e^{-x}-1}{x}$ are increasing, so that the left hand side of (5.35), as a function of $Q$, is decreasing when $a=1$. Also, it decreases from $\infty$ to -$\infty$, crossing the $Q$-line only once. By looking at the properties of the functions $\frac{e^{x}-1}{x}$ and $\frac{e^{-x}-1}{x}$ it is also possible to give more precise statements about the intervals on the $Q$-line containing the root. For example, if $C\in(0,C^{*})$ where $C^{*}=\frac{x^{*}}{2}$ and $x^{*}$ is the only real root of the equation $\frac{e^{x}-1}{x}=2$ ($x^{*}\sim 1.25643$), then $Q_{0}\in(C,x^{*}-C)$. In general, if $C$ is large and negative, then $Q_{0}\sim C$, whereas if $C$ is positive then $Q_{0}\in(0,x^{0})$, decreasing to 0 when $C$ increases. Here $x^{0}$ corresponds to the root of (5.35) when $C=0$, i.e. $x^{0}\sim 0.93082$.
This equilibrium position is stable and indeed is a center. In fact, the linearization of the dynamical system corresponding to the Hamiltonian (4.8) around $Q_{0}$ follows by letting $P=\delta p$ and $Q=Q_{0}+\delta q$ with $p$ and $q$ satisfying
$\dot{q}=p,\quad\dot{p}=G(Q_{0})q,$ (5.36)
where $G(Q)$ is the derivative of the left hand side of (5.35). We have shown however that the left hand side of (5.35) is decreasing, so the solution of (5.36) is oscillatory.

### v2

@@ 5.1 An example for the discrete case
$\begin{split}&\epsilon_{n}=\frac{1}{\sqrt{2N}}\tilde{\epsilon}_{n},\;\\ &\eta_{n}=\sqrt{\frac{e^{2a}-1}{2(1-e^{-2aN})}}e^{-a|n|}\tilde{\epsilon}_{n}\\ &k_{n}=Ce^{-a|n|}\end{split}$ (128)
where $\tilde{\epsilon}_{n}$ is the sequence given in (32). In the previous it is assumed that $\tilde{\epsilon}_{n}$ has compact support in [N,N][-N,N]. Both $\epsilon_{n}$ and $\eta_{n}$ are normalized, in the $\ell^{2}$ norm, to the unity.
From the general formulae (42) and (46) we get
$\dot{Q}=P,\quad\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (129)
Let us look at the sum defining $\dot{P}$. By taking the Taylor series of the exponential we can write
$\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\eta_{n}\frac{(Q\eta_{n}+\kappa_{n})^{k}}{k!}.$ (130)
@@
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\eta_{n}\frac{(Q\eta_{n}+\kappa_{n})^{k}}{k!}=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}$ (131)
@@
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}=\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{(1-\cos(\pi m))}{2}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}.$ (132)
We now recall, from (128), that when $m+1$ is even $\eta^{m+1}_{n}=w^{m+1}e^{-a(m+1)|n|}$, where $w=\sqrt{\frac{e^{2a}-1}{2}}$. By considering also the explicit formula for $\kappa_{n}$ from (128) and by performing the sum over $n$ we can write
$\sum_{n\in{\mathbb{Z}}}\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}\frac{Q^{m}\eta^{m+1}_{n}\kappa^{k-m}_{n}}{k!}=\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}(1-\cos(\pi m))\frac{Q^{m}w^{m+1}_{n}C^{k-m}}{k!(e^{a(k+1)}-1)}$ (133)
@@
$\begin{split}&\sum_{k=0}^{\infty}\sum_{m=0}^{k}{k\choose m}(1-\cos(\pi m))\frac{Q^{m}w^{m+1}_{n}C^{k-m}}{k!(e^{a(k+1)}-1)}=\\ &=w\sum_{k}\frac{(C+Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}-w\sum_{k}\frac{(C-Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}\end{split}$ (134)
@@
$\dot{P}=1-w\sum_{k}\frac{(C+Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}+w\sum_{k}\frac{(C-Qw)^{k}}{k!}\frac{1}{e^{a(k+1)}-1}$ (135)
We notice that the series (135) can be also expressed in terms of the following function:
$F(z)=\sum_{n=1}r^{n}z^{r^{n}}=z\frac{d}{dz}\sum_{n=1}(z^{r^{n}}-1),$ (136)
@@
$\begin{split}&\dot{Q}=P,\\ &\dot{P}=1+w(F(e^{C-Qw})-F(e^{C+Qw}))\end{split}$ (137)
where the function $F(z)$ defined in (136) satisfies the functional equation
$rF(z^{r})=F(z)-rz^{r},\quad r\doteq e^{-a},$ (138)
@@
Alternatively but equivalently, the equations of motion (137) can be defined by the function $f(z)=F(\exp(z))$, given by the series
$f(z)=\sum_{k=0}\frac{z^{k}}{k!}\frac{r^{k+1}}{1-r^{k+1}}.$ (139)
@@
$\sum_{n=1}^{N}r^{n}\exp(zr^{n}),$ (140)
by expanding the exponential in its Taylor series, summing over $n$ and letting $N\to\infty$ one gets $F(\exp(z))=f(z)$, with $f(z)$ defined as in (139).
@@ 5.1.1 The numerics of the discrete case
$f(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\frac{1}{3^{\frac{k+1}{2}}-1}.$ (146)
@@
$|R_{N}|=|\sum_{k=N+1}^{\infty}\frac{z^{k}}{k!}\frac{1}{3^{\frac{k+1}{2}}-1}|\leq\frac{\sqrt{3}}{2}\sum_{k=N+1}^{\infty}\frac{(\frac{|z|}{\sqrt{3}})^{k}}{k!}=\frac{\sqrt{3}}{2}\frac{(\frac{|z|}{\sqrt{3}})^{N+1}}{(N+1)!}\exp(\frac{\theta|z|}{\sqrt{3}}),$ (147)
where $\theta\in(0,1)$ and we used the Lagrange form of the reminder for the expansion of $\exp(\frac{|z|}{\sqrt{3}})$. Just to get an idea, for $N=30$ the previous formula gives an $R_{30}$ of the order $10^{-8}$ for $z=10$ and $\theta=1$ and an $R_{40}$ of $10^{-16}$ for $N=40$ and the same values of $z$ and $\theta$, whereas the values of the sums from $k=0$ to $k=30$ and from $k=0$ to $k=40$ are about $2\cdot 10^{2}$. We use the series up to $N=40$ and integrate numerically the equations of motion with a standard Runge-Kutta-Fehlberg method with Maple. The Figure (2) shows the numerical integration of the system (145) with three different initial conditions: $(Q(0),P(0))=(10,0)$, $(Q(0),P(0))=(8,5)$ and $(Q(0),P(0))=(5,0)$. The corresponding values of the energies are $607.38102$, $207.09954$ and $33.39845$. The plots on the right, in blue, are the plots of the analytical curves corresponding to the level curves defined by the Hamiltonian in these three cases and coincide with the orbits. The curve in red is the curve defined by $\dot{P}=0$: the equilibrium position is where this curve intersect the $Q$ axis.
Figure 2: Plot of three orbits of the system (145): on the left there is the numeric, on the right the plot of the corresponding Hamiltonian level curves. In red the nullclines defined by $\dot{P}=0$ and $\dot{Q}=0$.
@@
As specified in the Remark (3.8), the equations of motion (62) holds true also if the sequences $\eta_{n}$ and $\kappa_{n}$ belong to $\ell^{1}$. The properties of the orbits given in the Remark (3.6) or in the Remark (3.7) are also valid in these cases. As an example of the differential equations that one obtains, we can take $\epsilon_{n}$ like in (50) (hence $\kappa_{n}$ odd)
$\begin{split}&\eta_{n}=\frac{3\sigma}{\pi^{2}n^{2}},\;\;n\neq 0,\quad\eta_{0}=0,\quad\sigma^{2}=1\\ &\kappa_{n}=\frac{C\text{sgn}(n)}{n^{2}}\;\;n\neq 0,\quad\kappa_{0}=0.\end{split}$ (148)
From formula (62), to get $\dot{P}$ we need to evaluate the following sum
$\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}),$ (149)
and, by using explicitly the relations (148) we get
$\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\frac{3\sigma}{\pi^{2}}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\left(\exp\left(\frac{\frac{3\sigma}{\pi^{2}}Q+C}{n^{2}}\right)+\exp\left(\frac{\frac{3\sigma}{\pi^{2}}Q-C}{n^{2}}\right)\right).$ (150)
@@
$\zeta_{even}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{k!}\zeta(2k+2),$ (151)
@@
$\dot{Q}=P,\quad\dot{P}=1-\frac{3\sigma}{\pi^{2}}\zeta_{even}\left(\frac{\frac{3\sigma}{\pi^{2}}Q+C}{n^{2}}\right)-\frac{3\sigma}{\pi^{2}}\zeta_{even}\left(\frac{\frac{3\sigma}{\pi^{2}}Q-C}{n^{2}}\right)$ (152)
@@ 5.3 An example for the continuous case
$\begin{split}&\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\tilde{\epsilon}^{2}=1.\\ &\eta(x)=\frac{a}{\sqrt{1-\exp(-2a^{2}L)}}\exp(-a^{2}|x|)\tilde{\epsilon}(x),\\ &\kappa(x)=C\exp(-a^{2}|x|).\end{split}$ (153)
Let us look at the equations of motion for $P$ (71) and $Q$ (72). From the general formulae (103) and (108) we get
$\dot{Q}=P,\quad\dot{P}=1-\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi$ (154)
@@
$\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi=\sum_{n=0}^{\infty}\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}$ (155)
Since $\eta$ is an odd function, whereas $\kappa$ is even, only when $j$ is odd there is a contribution different from zero. So, by taking into account the expressions (153) and by noticing that $\tilde{\epsilon}^{j+1}=1$ for $j$ odd we can write:
$\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}=\sum_{j=0}^{n}{n\choose j}\frac{1-\cos(\pi j)}{2}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}C^{n-j}\alpha^{j+1}e^{-(n+1)\alpha^{2}|\xi|}}{n!}$ (156)
@@
$\sum_{j=0}^{n}{n\choose j}\int_{-\infty}^{\infty}d\xi\frac{Q^{j}\eta(\xi)^{j+1}\kappa(\xi)^{n-j}}{n!}=\sum_{j=0}^{n}{n\choose j}\frac{\alpha^{j-1}Q^{j}C^{n-j}(1-\cos(\pi j))}{(n+1)!}.$ (157)
The sum over $j$ can be carried out by using the formula for the binomial expansion. The sum in equation (157) is just given by
$\frac{(C+Qa)^{n}}{a(n+1)!}-\frac{(C-Qa)^{n}}{a(n+1)!}.$ (158)
Going back to equation (155) we can write
$\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi=\sum_{n=1}\left(\frac{(C+Qa)^{n}}{a(n+1)!}-\frac{(C-Qa)^{n}}{a(n+1)!}\right),$ (159)
@@
$\dot{Q}=P,\quad\dot{P}=1-\frac{e^{C+aQ}-1}{a(C+aQ)}+\frac{e^{C-aQ}-1}{a(C-aQ)}.$ (161)
@@
$\frac{e^{C+aQ}-1}{a(C+aQ)}+\frac{e^{C-aQ}-1}{a(C-aQ)}=1.$ (162)
As expected, this equation has always just one real root. For the sake of simplicity, let us just set $a=1$. Then, both the functions $\frac{e^{x}-1}{x}$ and $\frac{e^{-x}-1}{x}$ are increasing, so that the left hand side of (162), as a function of $Q$, is decreasing when $a=1$. Also, it decreases from $\infty$ to -$\infty$, crossing the $Q$-line only once. By looking at the properties of the functions $\frac{e^{x}-1}{x}$ and $\frac{e^{-x}-1}{x}$ it is also possible to give more precise statements about the intervals on the $Q$-line containing the root. For example, if $C\in(0,C^{*})$ where $C^{*}=\frac{x^{*}}{2}$ and $x^{*}$ is the only real root of the equation $\frac{e^{x}-1}{x}=2$ ($x^{*}\sim 1.25643$), then $Q_{0}\in(C,x^{*}-C)$. In general, if $C$ is large and negative, then $Q_{0}\sim C$, whereas if $C$ is positive then $Q_{0}\in(0,x^{0})$, decreasing to 0 when $C$ increases. Here $x^{0}$ corresponds to the root of (162) when $C=0$, i.e. $x^{0}\sim 0.93082$.
This equilibrium position is stable and indeed is a center. In fact, the linearization of the dynamical system corresponding to the Hamiltonian (75) around $Q_{0}$ follows by letting $P=\delta P$ and $Q=Q_{0}+\delta Q$ with $\delta P$ and $\delta Q$ satisfying
$\dot{\delta Q}=\delta P,\quad\dot{\delta P}=G(Q_{0})\delta Q,$ (163)
where $G(Q)$ is the derivative of the left hand side of (162). We have shown however that the left hand side of (162) is decreasing, so the solution of (163) is oscillatory.